OpenStudy (anonymous):

find zeros of 3x^2+2x-1

OpenStudy (anonymous):

3x^2 +3x - x -1 = 3x(x+1) -1(x+1) =(3x-1)(x+1) zero of its are -1 and 1/3 .

OpenStudy (anonymous):

thank you! now -x-2x+4

OpenStudy (anonymous):

it is -3x+ 4 so x= 4/3

OpenStudy (anonymous):

sorry it was -x^2-2x+4

OpenStudy (anonymous):

sorry, for late reply, something wrong with this site,....... can u use this link........ http://www.twiddla.com/493016